Circular No. 5707

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRB 930131 The Compton Observatory EGRET Team (M. Sommer, K. Brazier, G.

Kanbach, H. A. Mayer-Hasselwander and C. von Montigny, Max-Planck-Institut für Extraterrestrische Physik; D. L. Bertsch, B. L. Dingus, C. E. Fichtel, R. C. Hartman, S. D. Hunter, J. R. Mattox, P. Sreekumar and D. J. Thompson, NASA Goddard Space Flight Center; D. A. Kniffen, Hampden-Sydney College; J. Chiang, J. Fierro, Y. C. Lin, P. F. Michelson and P. L. Nolan, Stanford University; and E. Schneid, Grumman Aerospace Corporation), together with coinvestigator K. Hurley, University of California, report that the EGRET detector on the Gamma Ray Observatory imaged the Jan. 31 γ-ray burst at energies above 30 MeV. The highest event energies are significantly greater than have been observed in any previous burst. The location is α = 12^{h}11^{m}, δ = −10^{o}.7 (equinox 2000.0; 68-percent-confidence radius 40′; 95-percent-confidence radius 66′), near the edge of the error box reported on IAUC 5702; see also IAUC 5704.

RE J0751+14 V. Piirola and P. Hakala, University of Helsinki and Vatican Observa-

tory communicate: “We report the discovery of variable circular and linear polarization over the 13.9-min spin period of the intermediate polar RE J0751+14 (α = 7^{h}51^{m}17^{s}.3, δ = +14^{o}44′23′′, equinox 2000.0; V = 14.1) from observations with the Nordic 2.5-m telescope on 1992 Oct. 26. The total (peak-to-peak) circular polarization variation was 4.2 ± 0.6 percent in the I band (830 nm) and 2.2 ± 0.3 percent in the R band (690 nm), with slightly stronger negative extremes and a faster transition from the negative to the positive values. The linear polarization reached ∼ 2 percent in I and 0.8 percent in R. A preliminary analysis suggests cyclotron emission from two regions near the opposite magnetic poles of the white dwarf and field strength in the range 5-10 MG. The position-angle variability could be tentatively explained with an inclination of the magnetic white dwarf rotation axis 45^{o}-70^{o} and the two cyclotron-emission regions 35^{o}-60^{o} from the rotation poles. The intensity modulation was 0.18 mag in I and 0.10 mag in R, with a double-peaked structure, but in the UBV bands the lightcurve was nearly sinusoidal over the spin period, and the amplitude decreased to ∼ 0.07 mag. The UBV power spectra peak at 14.5 min, which could be the beat period corresponding to a binary period of ∼ 5 hr. Further polarimetric, spectroscopic and photometric observations are encouraged.”

1993 February 5 Brian G. Marsden